if sin1 sin2 sin3 …...sin89 =1/2^n. find the value of n. 

Grade:12th pass

1 Answers

Kashish
6 years ago
Sin1sin2sin3.........sin87sin88sin89=1/2^nSin(1+89)sin (2+88)sin (3+87).......sin (45+45)=1/2^nSin90+sin90........sin90=1/2^nTotal terms are 451/2^45=1/2^nn=45